OpenStudy (anonymous):

Find S∞ (if it exists)? 100-10+1-1/10+...

OpenStudy (anonymous):

\[\sum_{0}^{\infty}(10)^{-n}*(-1)^n\]

OpenStudy (anonymous):

\[\sum_{0}^{\infty}(((10))^{-1})^{n} * (-1)\] \[\sum_{0}^{\infty}({1 \over 10})^{n} * (-1)^{n}\] \[\sum_{0}^{\infty}({-1 \over 10})^{n} \]

OpenStudy (anonymous):

It is a geometric series

OpenStudy (anonymous):

so we use its formula 1/(1-r) 1/(1-(-1/10)) 1/(1+1/10)

OpenStudy (anonymous):

1/(11/10) 10/11